Fermi National Laboratory


Accelerator Update for February 14, 2005 - February 16, 2005

Monday February 14
The day shift began with Operations monitoring store 3980, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam, but with no beam to Meson experiments due to a closed Switchyard (SY) vacuum pump valve.

A vacuum expert got the SY valve at G2 open at 8:49 AM. Operations established beam to MCenter, but couldn't get beam to MTest.

Meson experts established beam to MTest at 10:25 AM.

At 11:23 AM, the Main Injector (MI) beam permit tripped off due to MECAR. It tripped off two more times within an hour. Operators called in an expert.

At 11:32 AM, Operations turned off Booster RF stations 8-18 so that RF technicians could repair and bring up BRF15. This halted stacking and all downstream beam.

Pbar resumed stacking at 12:38 PM.

NuMI experts began a study period at 2:23 PM.

At 2:34, Operations reported the loss of vacuum in the AP1 Pbar transfer line. Experts began their investigation. This halted stacking.

Operators prepared AP1 for access at 3:24 PM.

The MTest experiment asked to have their beam disabled at 3:56 PM.

Vacuum techs accessed the Pbar Pre-Vault area and hooked up another vacuum pump to the area with bad vacuum.

At 6:58 PM, Operations reported that a Booster RF station (BRF4) tripped off on a power supply overload. It reset okay.

At 7:10 PM, Operations reported that BRF15 had tripped off twice due to a spark. It reset okay.

At 8:49 PM, Operations reported that BRF17 had tripped off. It reset okay.

At 8:57 PM, Operations reported that BRF17 had tripped off for the first of eight times in about an hour. The last time it tripped was on a ground fault. Operators bypassed the station.

Tuesday February 15
The midnight shift began with Operations monitoring store 3980, the Recycler stash and the stack, with Pbar not stacking due to a vacuum problem, and with MiniBooNE and MCenter taking beam.

At 1:03 AM, Operators discovered that a number of MI vacuum pumps had tripped off on Monday. They were able to reset and turn on all but one of the pumps.

At 1:41 AM, Operators reported that they reset tripped vacuum pumps in the AP1 transfer line.

At 1:46 AM, Operations reported that vacuum valves closed in the MI RF section. Operator investigated but didn't find any pumps off, but they did find their permits off. Everything reset okay.

At 2:20 AM, Operations reported that MI suffered a RF trip due to a bad vacuum indication. It reset okay. After some investigation, operators discovered that a small vacuum burst had caused pressure spikes big enough to close the valves but not big enough to trip off the vacuum pumps.

Operations turned on BRF17 at 4 AM.

Operations terminated store 3980 at 12:06 PM.

At 1:37 PM, Operations reported that the Recycler stash was lost due to vacuum valves closing. Experts discovered that a vacuum pump that wasn't needed was turned on and caused a problem.

At 3:30 PM, Operations reported that an air compressor for Tevatron (TeV) sector F1 tripped off. An expert investigated and discovered that a Controls crate had failed. An expert replaced the crate power supply.

At 3:57 PM, BRF7 & 8 tripped off due to vacuum high pressure. They both reset okay.

At 5:03 PM, a Linac RF station (LRF5) tripped off. An expert investigated and discovered the station had tripped on a low nitrogen pressure permit.

At 5:05 PM, Operations reported that the TeV suffered a sector A13 quench due to DC beam in the abort gap.

Since the vacuum pressure had gotten better in AP1, a Pbar expert gave permission to resume stacking at 7:13 PM.

Pbar resumed stacking at 7:50 PM.

At 9:19 PM, BRF4 tripped off. It initially reset okay, but then it tripped off several more times. Operators bypassed the station.

Operations put the TeV through a dry squeeze at 10:14 PM. TeV experts resumed their studies about thirty minutes later.

Wednesday February 16
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE, MCenter, and MTest taking beam.

At 1:01 AM, a Pbar expert reported that a BPM was working properly in the AP1 line.

The MTest experiment asked to have their beam halted at 1:06 AM.

TeV experts completed their studies at 2:01 AM.

Operations put the TeV into a wet squeeze at 2:55 AM.

Operations put the TeV into shot setup at 3:16 AM.

Operations established store 3986 at 5:47 with an initial luminosity of 58.0E30. Pbar resumed stacking.

MCenter resumed taking beam at 6 AM.

At 6:37 AM, Operations reported that the MI beam permit had dropped out a few times since the shot due to a MECAR input.

The Plan for Wednesday
The plan for today is to monitor the store while Pbar and Booster are allowed to access their tunnels for required maintenance around 10 AM.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/16/2005   email Fermilab